# Gandalf the Red: Adaptive Security for LLMs

### v1

@@ 3.1  Analyzing D-SEC through Gandalf
We want to use Gandalf to investigate the security-utility trade-off in D-SEC and illustrate how to optimize defenses to maximize developer utility. We consider two data sources: Firstly, we created a modified and randomized version of Gandalf and collected a high-quality dataset, denoted by Gandalf-RCT (where RCT stands for Randomized Control Trial with the defense as treatment), which we use as attacker data to evaluate security, see Section 3.1.1. Secondly, we use a publicly available dataset BasicUser and a hand-crafted dataset BorderlineUser as benign user transactions to assess utility, see Section 3.1.2.
To ensure a diverse set of defenses, we created $6$ levels for which the application $\mathcal{M}$ is fixed, and we vary the defense $\mathcal{D}$. To investigate the effect of restricting the application domain, we implemented three setups representing different application domains: (1) The general setup representing an open-ended chatbot like ChatGPT, (2) the summarization setup representing an LLM summarization application where the attacks are more similar to indirect prompt attacks, and (3) the topic setup representing a narrow chatbot that is constrained to a specific topic as is common for customer support chatbots.
While there are some differences in the details, the high-level defenses implemented across the $6$ levels for each of the applications are as follows:
  • ATutorial: No defenses. The application $\mathcal{M}$ receives the password in its system prompt but is not explicitly instructed to avoid revealing the password.
@@
  • C1Substring check: Combines defense B with a check for the presence of certain substrings (e.g., “password” and the actual password itself) in the user message and application response. If either contains a forbidden string, the response is blocked, and the user is shown a pre-defined refusal message.
  • C2LLM checker: Combines defense B with a separate LLM call (same model as the application $\mathcal{M}$) to determine whether the user’s behavior is malicious. The checker decides based on both the user message and the application response. If the LLM checker flags the interaction, the response is blocked, and the user is shown a pre-defined refusal message.
@@
Full details of the defenses and setups, including the system prompts, can be found in Appendix A.
To draw generalizable (causal) conclusions from the data, we randomized levels over three axes: the defenses, underlying LLMs, and setups. More specifically, for each player (identified by a random session ID), we randomly selected a setup (general, summarization, or topic) and one of the following three OpenAI LLMs: GPT-3.5 (gpt-3.5-turbo-0125), GPT-4o-mini (gpt-4o-mini-2024-07-18) and GPT-4 (gpt-4-0125-preview). The selection of LLMs provides two axes of variation: (i) model size, which is believed to increase from GPT-3.5 and GPT-4o-mini to GPT-4 and (ii) model sophistication, which increases from GPT-3.5 and GPT-4 to GPT-4o-mini.
@@ 3.1.1 Attacker Data: Gandalf-RCT
The data collection for the randomized Gandalf experiment Gandalf-RCT was online from 2024-10-01 to 2024-11-07. In total, we collected 279’675 prompts in 36’239 sessions from 15’402 users across all setups and levels. We ran a detector of personally identifiable information (PII) on the data and filtered out prompts containing PII. Details on the data collection and supporting analyses not directly related the main message of this paper are provided in Appendices B and D. We additionally classified all attacks according to the triggers they use, similar as in Schulhoff et al. (2023), using active learning (Cohn et al., 1996). Details on all attack categories and the categorization procedure are provided in Appendix G, which illustrates the diversity of the attacks and provides interesting insights into different attack strategies.
@@ 3.1.2 User Data: BasicUser, BorderlineUser
Both datasets only contain single-transaction sessions. This is sufficient to evaluate non-adaptive defenses, however, to also be able to evaluate adaptive defenses in Section 4.2, we later artificially construct longer user sessions by independently sampling transactions from these datasets such that the user session length distribution matches that of the attack sessions.
@@ 4.1  Sensitivity of Utility to Data and Metrics
The resulting security and utility estimates for each defense (level-LLM combination) and for varying benign user data (BasicUser and BorderlineUser) are shown in Figure 3. For most defenses, there are large differences between the utility computed based on BasicUser and BorderlineUser, demonstrating the importance of choosing realistic user data when estimating utility (e.g., messages sent to an IT support bot might lead to a more borderline user distribution than that given by BasicUser). Our choice of user data is somewhat arbitrary as we are not considering an actual application, however in practice, the user pool is generally well defined and it is easy to collect representative user data.
As noted in Section 2, SCR captures utility loss only from blocked or refused requests. However, defenses integrated into the LLM, such as system prompt defenses, can subtly affect application behavior without outright blocking, impacting utility in less obvious ways. To illustrate this, we compare an undefended LLM (Level A), expected to maximize utility, with system prompt-defended variants (Levels B and C3). Using response character length and cosine similarity as proxies for utility, Figure 4 shows that strong defenses significantly reduce response length and alter content, suggesting degraded response quality. Even weaker defenses (Level B) exhibit measurable, though smaller, reductions. These results emphasize that system prompt defenses while enhancing security can subtly impact utility by changing model behavior beyond simple request blocking.
Figure 4: System prompt defenses degrade utility beyond simple request blocking. There is a noticeable impact of system prompt defenses on output quality, reflected in reduced response length (left) and altered content similarity (right). This suggests that even when user requests are not blocked, these defenses can still degrade utility. Triangles denote the mean, the box the 25%, 50% and 75% quantiles, respectively, and the whiskers the minimum and maximum values.
We stress that we did not optimize the defenses equally. For example, more work went into refining the strong system prompt than into setting up the LLM checker. One should, therefore, not use these results to draw general conclusions about the relative security of the different types of defenses (system prompt, LLM checker, and substring checks). We believe such a comparison is generally difficult and depends strongly on the setup.
@@ 4.2  Defense Strategies in D-SEC
Table 1: Average levels solved per session. Summarization is consistently the most difficult setup, followed by topic, suggesting that a more restricted use-case can improve defenses.
@@
The estimated AFR and SCR with BorderlineUser for adaptive defenses with varying block threshold $T$ based on each of the defenses in levels C1, C2, C3, and D (averaged over the LLMs) is shown in Figure 6. This shows that adaptive defenses with a low block threshold are able to significantly increase security over non-adaptive defenses that do not block entire sessions but only individual transactions. For instance, using a block threshold of 3 with the combined defense is able to block 75% of the attacks that were not caught by any of the non-adaptive attacks. The added security comes at the cost of utility, hence, in practice, we recommend to select the block threshold such that it maximizes developer utility. More detailed results are provided in Appendix F.2.

### v2

@@ 3.1  Analyzing D-SEC through Gandalf
We want to use Gandalf to investigate the security-utility trade-off in D-SEC and how to optimize defenses to maximize developer utility. We consider two data sources: Firstly, we created a modified and randomized version of Gandalf and collected a high-quality dataset, denoted by Gandalf-RCT (where RCT stands for Randomized Control Trial with the defense as treatment), which we use as attacker data to evaluate security, see Section 3.1.1. Secondly, we use a publicly available dataset BasicUser and a hand-crafted dataset BorderlineUser as benign user transactions to assess utility, see Section 3.1.2.
We define 6 levels where we keep the application $\mathcal{M}$ fixed while varying the defense $\mathcal{D}$. To study the effect of domain restriction, we implemented three setups: (1) The general setup representing an open-ended chatbot like ChatGPT, (2) the summarization setup representing an LLM summarization application with indirect prompt attacks, and (3) the topic setup representing a narrow chatbot focused on specific topics, as in customer support.
While there are differences in the details for different setups, the high-level defenses implemented across each of the $6$ levels are as follows (details are provided in Appendix A):
  • ATutorial: No defenses. The application $\mathcal{M}$ knows the password but is not told to avoid revealing it.
@@
  • C1Substring check: Combines defense B with a check for certain substrings (e.g., “password”) in user or application response. If the defense triggers, the response is blocked, and a pre-defined refusal message is shown.
  • C2LLM checker: Combines defense B with a separate LLM call (same LLM as for $\mathcal{M}$), using both the user and response message, to determine whether the user’s behavior is malicious. If the defense triggers, the response is blocked, and a pre-defined refusal message is shown.
@@
To draw generalizable (causal) conclusions, we randomize levels over three axes: defenses, LLMs, and setups. For each player (identified by a session ID), we randomly selected a setup (general, summarization, or topic) and one of three OpenAI LLMs: GPT-3.5 (gpt-3.5-turbo-0125), GPT-4o-mini (gpt-4o-mini-2024-07-18) and GPT-4 (gpt-4-0125-preview). These LLMs provide two axes of variation: model size, believed to increase from GPT-3.5 and GPT-4o-mini to GPT-4 and model sophistication, which increases from GPT-3.5 and GPT-4 to GPT-4o-mini.
@@ 3.1.1 Attacker Data: Gandalf-RCT
The data collection for the randomized Gandalf experiment Gandalf-RCT was online from 2024-10-01 to 2024-11-07. We collected 279’675 prompts in 36’286 sessions from 15’448 users across all setups and levels. We ran a detector of personally identifiable information (PII) on the data and removed prompts with PII. Further details on the data collection and supporting analyses are provided in Appendices B and D. We additionally classified all attacks according to their triggers as in Schulhoff et al. (2023), using active learning (Cohn et al., 1996). Details on all attack categories and the categorization procedure are provided in Appendix G, which illustrates attack diversity and provides insights into attack strategies.
@@ 3.1.2 User Data: BasicUser, BorderlineUser
Both datasets only contain single-transaction sessions. This is sufficient to evaluate non-adaptive defenses, however, to also be able to evaluate adaptive defenses in Section 4.2, we artificially construct longer sessions by independently sampling transactions from these datasets such that the distribution of session lengths equals that of attacker sessions.
@@ 4.1  Sensitivity of Utility to Data and Metrics
The resulting security and utility estimates for each defense (level-LLM combination) and for varying benign user data (BasicUser and BorderlineUser) are shown in Figure 3. For most defenses, there are large differences between the utility on BasicUser and BorderlineUser, demonstrating the importance of choosing realistic user data when estimating utility (e.g., messages sent to an IT support bot might lead to a more borderline distribution than that given by BasicUser). Our choice of user data is somewhat arbitrary as we are not considering an actual application, however in practice, the user pool is generally well defined and it is easy to collect representative user data.
As noted in Section 2, SCR captures utility loss only from blocked requests. However, defenses integrated into the LLM, such as system prompt defenses, can affect application behavior without blocking. To illustrate this, we compare an undefended LLM (Level A), expected to maximize utility, with system prompt-defended variants (Levels B and C3). Using response character length and cosine similarity as proxies for utility, Figure 4 shows that strong defenses significantly reduce response length and alter content, suggesting degraded response quality. Even weaker defenses (Level B) show smaller but measurable reductions. These results emphasize that system prompt defenses while enhancing security can subtly impact utility by changing model behavior beyond simple request blocking.
Figure 4: System prompt defenses degrade utility beyond simple request blocking. There is a noticeable impact of system prompt defenses on output quality, reflected in reduced response length (left) and altered content similarity (right). This suggests that even when user requests are not blocked, defenses can still degrade utility. Triangles denote the mean, the box the 25%, 50% and 75% quantiles, respectively, and the whiskers the minimum and maximum values.
We stress that we did not optimize defenses equally. For example, more work went into setting up the strong system prompt than the LLM checker. Thus, these results should not be used to draw general conclusions about the relative security of the different types of defenses (system prompt, LLM checker, and substring checks). Such comparison are generally difficult and depend strongly on the setup.
@@ 4.2  Defense Strategies in D-SEC
Table 1: Average levels solved per session. Summarization is the most difficult setup, followed by topic, suggesting that a more restricted use-case can improve defenses.
@@
The estimated AFR and SCR with BorderlineUser for adaptive defenses with varying block threshold $T$ based on each of the defenses in levels C1, C2, C3, and D (averaged over the LLMs) is shown in Figure 6. This shows that adaptive defenses with a low block threshold are able to significantly increase security over non-adaptive defenses that block individual transactions only. For instance, using a block threshold of 3 with the combined defense is able to block 75% of the attacks that were not caught by any of the non-adaptive attacks. The added security comes at the cost of utility, hence, in practice, we recommend to select the block threshold such that it maximizes developer utility. More detailed results are provided in Appendix F.2.